Two tests only. Prof. Davis counts his midterm as 2/3 of the grade, and his final as 1/3 of the grade. Jason scored only 64 on the midterm. What range of scores on the final exam would put Jason's average between 70 and 79 inclusive?
Please kindly help me.

Answer by stanbon(75887)

Two tests only. Prof. Davis counts his midterm as 2/3 of the grade, and his final as 1/3 of the grade. Jason scored only 64 on the midterm. What range of scores on the final exam would put Jason's average between 70 and 79 inclusive?
-----------------------
Let the final grade need be "x"
INEQUALITY:
70 <= (2/3)64+(1/3)x <= 79
Subtract (2/3)64 = 128/3 along the line to get:
82/3 <= (1/3)x <= 109/3
Multiply thru by 3 to get:
82 <= x <= 109
==================
Cheers,
Stan H.